given f(x)5x^3, find f^-1(x). then state whether f^-1(x) is a function
Answer by stanbon(75887)

given f(x)= 5x^3,
find f^-1(x).
1st: Interchange x and y to get:
x = 5y^3
---
2nd: solve for "y":
y^3= (x/5)
---
y = cbrt(x/5)
-------------------
then state whether f^-1(x) is a function
Yes, it is a function.
---------------------
Cheers,
Stan H.